# A Resolution Independent Neural Operator

### v2

@@ Appendix F Ablation Study: Comparison with Gappy Proper Orthogonal Decomposition and Random Projection
As we have shown, linear (with respect to the basis functions) projection and reconstruction using linearly independent random basis functions often requires exponentially more basis functions compared to an optimal orthogonal basis set, provided the signals of interest can be spanned by such an optimal orthogonal basis. Although this is aligned with the linear theory [16], it does not take advantage of an additional possibility, consistent with established embedding theorems, which suggests that one can embed an $n$-dimensional manifold with at most $12n+$ 1 generic observables [43, 44] (in our case, the components of our data in $12n+$ 1 random basis functions). These projections do not span our intrinsically $n$-dimesional data; yet, the projections of the data in any additional random basis functions can be obtained as functions of the first $12n+$ 1 ones. We say that the $12n+$ 1 basis functions parametrize the data, but do not span them. The approximation of the data can now be achieved via nonlinear parameterization, which maps the projected solution on the hyperplane back to the target manifold. This approach shares similarities with the post-processing Galerkin method [20, 14]. However, such nonlinear parameterization may have some caveats: (1) it is less interpretable than our proposed spanning scheme; (2) the number of basis functions is not known a priori—only an upper bound is provided—while our approach adaptively identifies the number of basis functions according to a tolerance; (3) The embedding space constructed via random projection is highly sensitive to sampling, which undermines the uniqueness of the signal representation. To elaborate further on the last point, we conducted a simple test as follows. We pick one of the signal realization from the test data set and aim to reconstruct the signal with three linearly independent random cosine basis, as mentioned earlier. To assess the sensitivity of the embeddings, we consider five cases where half of the points are masked in each case. However, the locations of these masked points are randomly selected and vary from one case to another. The embeddings obtained from random projection, shown in the right column of Figure 32(a), are clearly dependent on the location of the samples. The left column shows the reconstructed signal, where we again observe high variance. In contrast, the projection via optimal orthogonal basis functions, as shown in Figure 32(b),is consistent across cases and independent of the sample locations.
@@
Figure 32: Consistency check of the embeddings obtained from projections via (a) random basis and (b) optimal basis sets. The right column display the coefficients corresponding to each basis function, while the left column show the reconstructed full fields based on the identified coefficients. Different colors represent distinct queried datasets, where the number of data points remains the same but their locations are chosen randomly and differ from one dataset to another.
Appendix G Random Field Generation

### v3

@@ Appendix F Ablation Study: Comparison with Gappy Proper Orthogonal Decomposition and Random Projection
As we have shown, arbitrary linear (with respect to the basis functions) projection and reconstruction using linearly independent random basis functions often requires exponentially more basis functions than projections using an optimal orthogonal basis set. Although this is aligned with the linear theory [19], it does not take advantage of the fact that, with established embedding theorems, one can embed an $n$-dimensional manifold with at most $12n+$ 1 generic observables [61, 62] (in the present case, the data components result in $12n+$ 1 random basis functions). These projections do not span the intrinsically $n$-dimesional data; yet, the projections of the data in any additional random basis functions can be obtained as (potentially nonlinear) functions of the first $12n+$ 1 random basis functions. We say that the $12n+$ 1 basis functions parametrize the data, but do not span them. The approximation of the data can now be achieved via nonlinear parameterization, which maps the projected solution on the hyperplane back to the target manifold. This approach shares similarities with the post-processing Galerkin method [30, 17]. However, such nonlinear parameterization may have some caveats: (1) it is less interpretable than the proposed spanning scheme; (2) the number of basis functions is not known a priori—only an upper bound is provided—while the proposed approach adaptively identifies the number of basis functions according to a tolerance; (3) The embedding space constructed via random projection is highly sensitive to sampling, which undermines the uniqueness of the signal representation. To elaborate further on the last point, we conducted a simple test as follows. We pick one signal realization from the test data set and aim to reconstruct the signal with three linearly independent random cosine basis functions, as mentioned earlier. To assess the sensitivity of the embeddings, we consider five cases where half of the points are masked in each case. However, the locations of these masked points are randomly selected and vary from one case to another. The embeddings obtained from random projection, shown by the coefficients in the right pane of Figure 35(a), are clearly dependent on the location of the samples. The left column shows the reconstructed signal (solid lines with different colors), where we observe high variance due to this sample location dependence. In contrast, the projection via optimal orthogonal basis functions, as shown in Figure 35(b),is consistent across cases and independent of the sample locations.
@@
Figure 35: Consistency check of the embeddings obtained from projections via (a) random basis and (b) optimal basis sets. The right column display the coefficients corresponding to each basis function, while the left column show the reconstructed full fields using the identified coefficients. Different colors represent distinct training datasets, where the number of data points remains the same but their locations are chosen randomly and differ from one dataset to another. For illustration purposes, the data from the first two of the five trials is shown in the top-left panel using black and blue “x” symbols, with their corresponding reconstructions represented by black and blue solid lines, respectively.
Appendix G Ablation Study: Comparison between RI-DeepONet and RINO
Here, we aim to demonstrate that while RI-DeepONet or DeepONet do not necessarily identify the most compact or orthogonal set of basis functions to span the output field, RINO effectively achieves this. The example used here is the antiderivative dataset introduced by Lu Lu et. al [34].
Since the focus is on the impact of output basis functions on prediction quality, no data is masked in this study. Instead, full, clean data is utilized across all cases. The same input embedding is employed for all scenarios, determined through the proposed dictionary learning approach applied to the input function data.
RINO identifies that only five output basis functions are sufficient to span the output functions effectively. Each of these basis functions is implemented as a SIREN with a single hidden layer, 10 hidden units, and a frequency parameter of $w_{0}=5$.A key advantage of RINO is its ability not only to identify these basis functions but also to determine the number required—an unknown hyperparameter in both DeepONet and RI-DeepONet. In this case, a linear transformation between the input and output embeddings proved sufficient and was determined analytically. The strong generalization capability of RINO is evident from the empirical cumulative distribution functions (ECDFs) of training and test relative errors, plotted on a logarithmic scale in Figure 36.
Figure 36: RINO ablation: (left) ecdf plots for the learned operator, (right) the learned output basis functions.
In the case of RI-DeepONet experiments, the branch network is a vanilla MLP with a single hidden layer of size 50 and an ELU activation function. We consider two types of trunk networks: one is a vanilla MLP with a single hidden layer of size 50, and the other is a SIREN with a single hidden layer of size 50 and a frequency parameter of 5. Since RINO has already established that the output function space can be effectively spanned by only five basis functions, we test three cases where the sizes of the last layers of the branch and trunk networks are set to 5, 10, and 20.
The ECDF plots for these RI-DeepONet six cases are presented in Figure 37. Comparing these ECDFs with the RINO ECDF, one can observe two key points. First, the generalization gap in RINO is consistently much tighter than in RI-DeepONet, as evidenced by the closer alignment of training and test ECDFs in RINO. Second, the worst-case test prediction error in RINO is consistently better than in RI-DeepONet. We hypothesize that this improvement can be attributed to the superior quality of the output basis functions learned by RINO. These basis functions align closely with the data behavior and, importantly, are the most compact, containing no unnecessary information due to their orthogonality. A similar generalization effect, attributed to identifying the correct basis functions, was also observed in Appendix F, though from the perspective of input functions.
(a) Vanilla MLP 5
(b) Vanilla MLP 10
(c) Vanilla MLP 20
(d) SIREN 5
(e) SIREN 10
(f) SIREN 20
Figure 37: RI-DeepONet ablation: ECDF plots for six different architectures. The branch net remains the same across all cases, while the trunk net differs in the number of hidden units (i.e., the number of output basis functions) and its architecture. The top row uses an MLP as the trunk net, and the bottom row uses a SIREN. Numbers 5, 10, and 20 indicate the number of hidden units in the trunk net.
To further examine the aforementioned hypothesis, we plot the trunk basis functions learned by RI-DeepONet in Figure 38. It is evident from these plots that the learned basis functions do not guarantee compactness or orthogonality. As a result, some basis functions may exhibit linear dependence or high correlation, introducing redundancy in the approximation. This redundancy reduces the efficiency of the representation and may lead to overfitting, as the model focuses on memorizing training data instead of accurately capturing the underlying data structure.
(a) Vanilla MLP 5
(b) Vanilla MLP 10
(c) Vanilla MLP 20
(d) SIREN 5
(e) SIREN 10
(f) SIREN 20
Figure 38: RI-DeepONet ablation: learned trunk basis functions for six different architectures. The branch net remains the same across all cases, while the trunk net differs in the number of hidden units (i.e., the number of output basis functions) and its architecture. The top row uses an MLP as the trunk net, and the bottom row uses a SIREN. Numbers 5, 10, and 20 indicate the number of hidden units in the trunk net.
The operator learning loss functions for RI-DeepONet are shown in Figure 39 to ensure that the training has not been extended unnecessarily, which could inadvertently lead to overfitting. As the results indicate, the training process was healthy, with no significant evidence of overfitting observed in the training data. All these observations support our hypothesis that the structure of the learned basis functions may influence accuracy, robustness, generalization, and interpretability. Consequently, algorithms that ensure the learning of well-structured basis functions with such guarantees are desirable.
(a) Vanilla MLP 5
(b) Vanilla MLP 10
(c) Vanilla MLP 20
(d) SIREN 5
(e) SIREN 10
(f) SIREN 20
Figure 39: RI-DeepONet ablation: operator learning loss functions for six different architectures. The branch net remains the same across all cases, while the trunk net differs in the number of hidden units (i.e., the number of output basis functions) and its architecture. The top row uses an MLP as the trunk net, and the bottom row uses a SIREN. Numbers 5, 10, and 20 indicate the number of hidden units in the trunk net.
Appendix H Random Field Generation